# Taylor-Fourier integration

### v1

@@ 3.1 Step 1: Evaluation of trigonometric polynomials
${\bf Y}_{n,j}={\bf Y}_{j}(\theta_{n})=\sum_{k=-M}^{M}e^{ikn\pi/M}{\bf y}_{k,j}^{[d]},\quad j=0,1,\ldots,d,$ (17)
by applying the inverse DFT
@@
$\begin{array}[]{l}\hat{\bf Y}_{0,j}=2M\,{\bf y}_{0,j}^{[d]},\quad\hat{\bf Y}_{M,J}=2M({\bf y}_{M,j}^{[d]}+{\bf y}_{-M,j}^{[d]}),\\[8.0pt] \hat{\bf Y}_{k,j}=2M\,{\bf y}_{k,j}^{[d]},\quad\hat{\bf Y}_{2M-k,j}=2M{\bf y}_{-k,j}^{[d]},\quad k=1,\ldots,M-1.\end{array}$
@@ 3.2 Step 2: Truncated Taylor expansions
${\bf f}\left({\theta_{n},}\sum_{j=0}^{d}t^{j}\,{\bf Y}_{n,j}\right)=\sum_{j=0}^{d}t^{j}\,{\bf Z}_{n,j}+{\cal O}(t^{d+1}).$
This means that we obtain the Taylor polynomial of degree $d$ of the left hand side of the previous equality. The computation of the required Taylor expansion can be done in a very efficient way when appropriate tools to perform power series arithmetic are available.
@@ 3.4 Step 4: Computing quadratures
To simplify the notation, we will drop the superscript $[d+1]$ from the vector coefficients ${\bf y}_{k,j}^{[d+1]}$ in the remaining of the present subsection, so that we will write
${\bf y}^{[d+1]}(t)=t^{d+1}\,{\bf y}_{0,d+1}+\sum_{k=-M}^{M}e^{ik\omega t}\left(\sum_{j=0}^{d}t^{j}\,{\bf y}_{k,j}\right).$
@@ 4 Numerical illustrations
In this section we include two numerical examples to illustrate the performance of the Taylor-Fourier approximations proposed in this paper, when compared with state-of-the-art methods. All codes we have used to carry out the numerical experiments are available in https://github.com/jmakazaga/Taylor-Fourier.
@@ 4.1 The cubic nonlinear Schrödinger equation
Products of a vector in $\mathbb{C}^{2J}$ by matrices $e^{\theta A}$ and $e^{-\theta A}$ can be computed efficiently by performing two FFT in $\mathbb{C}^{2J}$, since $e^{\theta A}={\cal F}^{-1}e^{\theta D}{\cal F}$, and $e^{\theta D}$ is the diagonal matrix with diagonal entries $(e^{\lambda_{1}\theta},\ldots,e^{\lambda_{2J}\theta})$. On the other hand, due to the particular form of the nonlinearity ${\bf g}$ in (30), the Taylor expansions required to apply the Taylor-Fourier integrator are cheaply available since sums and products of power series in $t$ can be computed in a very efficient way.
@@
Secondly, to illustrate how the choice of $d$ and $M$ affects the numerical solution computed with the Taylor-Fourier method, we include in Figure 2 the numerical approximations to the solution of (26)-(32) at $t=\pi/10$ when $J=2^{9}$, and the combinations $M=4J$, $d=1$ (left), $M=J$, $d=3$ (center) and $M=J$, $d=1$ (right) are used. Plots in Figure 2 must be compared with the bottom right plot in Figure 1 where the $(4J,3)$-Taylor-Fourier approximation is displayed. Comparing the plots corresponding to the same value of $M$ and reducing $d$ from 3 to 1, we observe that, although the overall shape of the numerical solutions is rather similar, the maximum and minimum values of both solutions are clearly different. Looking now at the solutions obtained with the same value of $d$ we see that when decreasing $M$ from $4J$ to $J$ small oscillations appear.
@@
Finally, we have tested the efficiency (time-discretization error $|\tilde{U}_{j}(t)-U_{j}(t)|$ versus CPU time) of Taylor-Fourier approximations to the solution $u(t,x)$ of (26) with initial data (33). We have considered different space discretization parameters depending on the value of $\epsilon$: $J=2^{9}$ for $\epsilon=1$, and $J=2^{6}$ for $\epsilon=2^{-4}$. We have plotted in Figure 4 the results corresponding to $(M,d)$-Taylor-Fourier integrations with different choices of the parameters $M$ and $d$.
The reference solution of the space-discretized problem used to measure time-discretization errors has been computed as indicated above. For comparison purposes, we have also included errors obtained with the same scheme for tolerances $10^{-3}$, $\ldots$, $10^{-6}$ when $\epsilon=1$, and for tolerances $10^{-3}$, $\ldots$, $10^{-7}$ for $\epsilon=2^{-4}$. In both plots, a black dashed line representing the spatial error for the chosen value of $J$ (estimated by a precise time-integration of the semi-discretized problem with $2J$ spatial nodes) has been included. Notice that although $J$ is larger for $\epsilon=1$ than for $\epsilon=2^{-4}$, the spatial error is also larger for $\epsilon=1$ than for $\epsilon=2^{-4}$. This is due to the factor $\epsilon$ that multiplies the rough initial data (32). In both cases we have used $2\leq d\leq 5$.
Figure 4: Errors against CPU time when $J=2^{9}$, the initial data is (32) and $T=\pi/10$ (left), and when $J=2^{6}$, the initial data is (33) with $\epsilon=2^{-4}$ and $T=\epsilon^{-2}\pi/10$ (right). Diamonds correspond to the Lawson method based on DP5 and circles to the $(M,d)$-Taylor-Fourier approximations
We observe in the left plot of Figure 4 that the Taylor-Fourier integrations are considerably less efficient than the Lawson-DP5 integrations for the case $\epsilon=1$. In the right plot, for $\epsilon=2^{-4}$, we see that the performance of both integrators is rather similar. The low efficiency of the Taylor-Fourier methods for $\epsilon=1$ is due to the high dimension of the semi-discretized problem ($J=2^{9}$), and to the fact that in order to get temporal errors not larger than the spatial errors, the parameter $M$ varies from $2^{8}$ to $2^{11}$, increasing the computational cost of the method. For $\epsilon=2^{-4}$, the dimension of the semi-discretized problem has been divided by $8$, and the problem is more difficult to be solved using standard integrators because its solution is really highly oscillatory. Then, although $M$ ranges between $2^{6}$ and $2^{9}$, the advantage of Lawson-DP5 with respect to Taylor-Fourier is much smaller.
@@ 4.2.2 Stiefel and Scheifele’s formulation
${\bf u}(\tau)=\cos(\omega\tau)\bm{\alpha}(\tau)+\omega^{-1}\sin(\omega\tau)\bm{\beta}(\tau),$
where $\bm{\alpha}(\tau)$ and $\bm{\beta}(\tau)$ are the solutions of the 8-dimensional ODE system
$\begin{split}\bm{\alpha}^{\prime}&=\omega^{-1}\sin(\omega\tau)\nabla R(\cos(\omega\tau)\bm{\alpha}+\omega^{-1}\sin(\omega\tau)\bm{\beta}),\\ \bm{\beta}^{\prime}&=-\cos(\omega\tau)\nabla R(\cos(\omega\tau)\bm{\alpha}+\omega^{-1}\sin(\omega\tau)\bm{\beta}),\end{split}$ (44)
@@
$\bm{\alpha}(0)={\bf u}_{0},\quad\bm{\beta}(0)={\bf u}^{\prime}_{0}.$ (45)
@@
$\bm{\alpha}(\tau)\approx\sum_{k=-M}^{M}\sum_{j=0}^{d}e^{ik\omega\tau}\,\tau^{j}\,\bm{\alpha}_{k,j},\quad\bm{\beta}(\tau)\approx\sum_{k=-M}^{M}\sum_{j=0}^{d}e^{ik\omega\tau}\,\tau^{j}\,\bm{\beta}_{k,j},$
of $\bm{\alpha}(\tau)$ and $\bm{\beta}(\tau)$ are obtained.

### v2

@@ 3.1 Step 1: Evaluation of trigonometric polynomials
${\bf Y}_{n,j}={\bf Y}_{j}(\theta_{n})=\sum_{k=-M}^{M}e^{ikn\pi/M}{\bf y}_{k,j},\quad j=0,1,\ldots,d,$ (26)
(with the nodes $\theta_{n}$ considered in (23)) by applying the inverse DFT
@@
$\begin{array}[]{l}\hat{\bf Y}_{0,j}=2M\,{\bf y}_{0,j},\quad\hat{\bf Y}_{M,j}=2M({\bf y}_{M,j}+{\bf y}_{-M,j}),\\[8.0pt] \hat{\bf Y}_{k,j}=2M\,{\bf y}_{k,j},\quad\hat{\bf Y}_{2M-k,j}=2M{\bf y}_{-k,j},\quad k=1,\ldots,M-1.\end{array}$
@@ 3.2 Step 2: Truncated Taylor expansions
${\bf f}\left({\theta_{n},}\sum_{j=0}^{d}t^{j}\,{\bf Y}_{n,j}\right)=\sum_{j=0}^{d}t^{j}\,{\bf Z}_{n,j}+{\cal O}(t^{d+1}).$ (27)
The required truncated Taylor expansion (27) can be computed efficiently without resorting to symbolic calculations by evaluating the univariate function on the left-hand side of (27) using truncated power series arithmetic. To achieve this, one can implement the mapping ${\bf f}(\theta,{\bf y})$ as a function that takes an angle $\theta$ and $D$ polynomials in $t$ of degree $d$, returning the corresponding $D$ polynomials of degree $d$. This approach has been used in both examples presented in Section 6. Alternatively, numerous libraries are available across different computing environments that support arithmetic for truncated power series.
@@ 3.4 Step 4: Computing quadratures
4 Taylor-Fourier methods and high order stroboscopic averaging
In this section, we establish a connection between semi-discrete Taylor-Fourier expansions and high-order stroboscopic averaging of systems of the form (9) based on the results in [4]. We begin by recalling from [4, 5] the series expansion (10) and its relation to high-order averaging, which, as shown in [4], is derived through successive applications of the Picard-like iteration (12)–(14). Note that, by rescaling time, we may assume without loss of generality that either $\omega=1$ or $\epsilon=1$ in (9). However, we retain this redundancy in this section, as it facilitates the extension of our results to the quasi-periodic case considered in Subsection 5.2.
In $d$th-order averaging, a change of variables of the form
${\bf y}={\bf U}(\omega t,{\bf w};\epsilon)={\bf w}+\epsilon\,{\bf u}_{1}(\omega t,{\bf w})+\cdots+\epsilon^{d}\,{\bf u}_{d}(\omega t,{\bf w})$
is introduced. This transformation converts the original system (9) into an averaged system of the form
$\frac{d}{dt}{\bf w}=\epsilon\,{\bf F}({\bf w};\epsilon)+\epsilon^{d+1}{\bf R}(\omega t,{\bf w};\epsilon),\quad{\bf w}(0)={\bf w}_{0},$
where ${\bf F}({\bf w};\epsilon)$ represents the averaged dynamics, and ${\bf R}(\omega t,{\bf w};\epsilon)$ captures the remaining small amplitude oscillatory terms.
In the stroboscopic averaging framework, the solution ${\bf y}(t)$ of (9) admits the representation
${\bf y}(t)={\bf U}(\omega t,{\bf W}(t,{\bf y}_{0};\epsilon);\epsilon)+\mathcal{O}(\epsilon^{d+1}),$ (36)
where ${\bf W}(t,{\bf y}_{0};\epsilon)={\bf w}(t)$ denotes the solution of the averaged system
$\frac{d}{dt}{\bf w}=\epsilon\,{\bf F}({\bf w};\epsilon),\quad{\bf w}(0)={\bf y}_{0}.$ (37)
Consequently, the solution ${\bf x}(t)$ of the original semi-linear equation
$\frac{d}{dt}{\bf x}=\mathcal{A}{\bf x}+\epsilon\,{\bf g}({\bf x})$ (38)
satisfies
${\bf x}(t)=e^{t\mathcal{A}}\,{\bf U}(\omega t,{\bf W}(t,{\bf y}_{0};\epsilon);\epsilon)+\mathcal{O}(\epsilon^{d+1}).$
As shown in [4, Theorem 2.10] (see also [5, Theorem 2.1]), we have that, for arbitrary ${\bf y}_{0}\in\mathbb{R}^{D}$,
$\begin{split}{\bf U}(\omega t,{\bf y}_{0};\epsilon)&={\bf y}_{0}+\sum_{\ell=1}^{d}\epsilon^{\ell}\sum_{k=-\infty}^{\infty}e^{ik\omega t}{\bf y}_{k,\ell,0},\\ {\bf W}(t,{\bf y}_{0};\epsilon)&={\bf y}_{0}+\sum_{\ell=1}^{d}\epsilon^{\ell}\sum_{j=0}^{\ell}t^{j}\sum_{k=-\infty}^{\infty}{\bf y}_{k,\ell,j},\\ {\bf F}({\bf y}_{0};\epsilon)&=\frac{\partial}{\partial t}{\bf W}(0,{\bf y}_{0};\epsilon)=\sum_{\ell=1}^{d}\epsilon^{\ell-1}\sum_{k=-\infty}^{\infty}{\bf y}_{k,\ell,1},\end{split}$ (39)
where ${\bf y}_{k,\ell,j}$ are the coefficients of the expansion (10) of the solution of (9). Recall that these coefficients can be obtained through successive applications of the Picard-like iteration (12)–(14).
The $(\infty,d)$-Taylor-Fourier approximation of the solutions ${\bf y}(t)$ and ${\bf x}(t)$ of (9) and (38) read
${\bf Y}^{[d]}(\omega t,t;\epsilon,{\bf y}_{0})=\sum_{j=0}^{d}\epsilon^{j}t^{j}\,\sum_{k=-\infty}^{\infty}e^{ik\omega t}{\bf y}_{k,j}^{[d]}.$ (40)
and ${\bf X}^{[d]}(\omega t,t;\epsilon,{\bf y}_{0})=e^{t\,\mathcal{A}}\,{\bf Y}^{[d]}(\omega t,t;\epsilon,{\bf y}_{0})$, respectively.
Expansions (40) and (11) are derived through a similar Picard-like iteration process. The key distinction lies in the truncation method used to approximate ${\bf z}(t)={\bf f}(\omega t,{\bf y}(t))$: (40) is obtained by truncating higher powers of $t$, whereas (11) involves truncating higher powers of $\epsilon$. This latter approach discards more terms in the expansion of ${\bf z}^{[d]}(t)={\bf f}(\omega t,{\bf y}^{[d]}(t))$. This implies that the coefficients of the expansions (40) and (11) are related by
$\epsilon^{j}{\bf y}_{k,j}^{[d]}=\epsilon^{j}{\bf y}_{k,j,j}+\mathcal{O}(\epsilon^{d+1}).$
We thus have the following relationships:
${\bf U}(\omega t,{\bf y}_{0};\epsilon)$ $={\bf Y}^{[d]}(\omega t,0;\epsilon,{\bf y}_{0})+\mathcal{O}(\epsilon^{d+1}),$
${\bf W}(t,{\bf y}_{0};\epsilon)$ $={\bf Y}^{[d]}(0,t;\epsilon,{\bf y}_{0})+\mathcal{O}(\epsilon^{d+1}),$
$\epsilon\,{\bf F}({\bf y}_{0};\epsilon)$ $=\left.\frac{\partial}{\partial t}{\bf Y}^{[d]}(0,t;\epsilon,{\bf y}_{0})\right|_{t=0}+\mathcal{O}(\epsilon^{d+1}).$
This demonstrates that the change of variables in the $d$th-order stroboscopic averaging can be effectively approximated by applying the $(M,d)$-Taylor-Fourier algorithm with a sufficiently large $M$. In particular, this together with (36) implies that
${\bf X}^{[d]}\left(\omega t,0;\epsilon,{\bf Y}^{[d]}(0,t;\epsilon,{\bf y}_{0})\right)={\bf X}^{[d]}(\omega t,t;\epsilon,{\bf y}_{0})+\mathcal{O}(\epsilon^{d}).$ (41)
While the right-hand side of the averaged equations and their solutions can also be computed using this approach, we do not claim it to be an efficient method for this purpose. However, it can serve as a useful tool to assess the accuracy of more efficient implementations of approximate averaged equations and their solutions.
We note that the semi-discrete expansions (11) in powers of $\epsilon$ can be transformed into a practical fully discrete method by adapting the procedure proposed in this paper. Specifically, for a given pair $(M,d)$, a finite-sum substitute for the $d$th order expansion (11) can be constructed by applying trigonometric interpolation to replace the periodic functions
$\sum_{k=-\infty}^{\infty}e^{ik\omega t}{\bf z}_{k,\ell,j}$
in the first sub-step of the iteration defined by (12)–(14).
This approach is particularly relevant in the context of stroboscopic averaging, especially when one aims to evaluate the mappings in (39) exactly –or with high precision for a sufficiently large $M$– for a given order $d$. In particular, the right-hand side ${\bf F}({\bf y}_{0};\epsilon)$ of the averaged equations obtained through this method preserves the geometric structure of the original equations. For instance, if the original equations (38) are Hamiltonian, then the averaged equations retain this Hamiltonian structure. In contrast, the approximation
$\left.\frac{\partial}{\partial t}{\bf Y}^{[M,d]}(0,t;\epsilon,{\bf y}0)\right|_{t=0}\approx{\bf F}({\bf y}_{0};\epsilon),$
obtained using the Taylor-Fourier algorithm proposed in Section 3, generally does not correspond to a Hamiltonian vector field. In this context, the truncated expansions (11) and (40) are not symplectic, as is expected for any explicit formula derived by truncating a power series expansion of a symplectic map.
However, this fully discrete method is computationally more expensive than the Taylor-Fourier algorithm proposed in this paper, both in terms of memory and complexity. For a given order $d$, it requires $(d+2)/2$ times more coefficients, and the algorithm involves arithmetic with truncated power series in two variables (rather than one).
5 Application to more general equations
5.1 Semi-linear equations in the complex domain
5.2 Semi-linear equations with non-resonant frequencies
Taylor-Fourier approximations can be naturally extended to systems of the form (1), where $\mathcal{A}$ is a real $D\times D$ matrix with purely imaginary eigenvalues that are not fully resonant. Applying the time-dependent change of variables
${\bf x}(t)=e^{t\mathcal{A}}{\bf y}(t),$ (42)
we transform (1) into
$\frac{d}{dt}{\bf y}=e^{-t\mathcal{A}}{\bf g}(e^{t\mathcal{A}}{\bf y}).$
In this general setting, there exists a vector of non-resonant frequencies $\boldsymbol{\omega}=(\omega_{1},\ldots,\omega_{r})\in\mathbb{R}^{r}$ such that $\mathbf{k}\cdot\boldsymbol{\omega}=k_{1}\omega_{1}+\cdots+k_{r}\omega_{r}\neq 0$ for all $\mathbf{k}=(k_{1},\ldots,k_{r})\in\mathbb{Z}^{r}/\{{\bf 0}\}$. The eigenvalues of $\mathcal{A}$ then take the form $(\mathbf{k}\cdot\boldsymbol{\omega})i$ for $\mathbf{k}\in\mathbb{Z}^{r}$, making the solution operator $e^{t\mathcal{A}}$ of the linear part of (1) quasi-periodic in $t$, rather than strictly periodic as in the fully resonant case. This implies that there exists a mapping
$\begin{split}{\bf f}:&\mathbb{T}^{r}\times\mathbb{R}^{D}\longrightarrow\mathbb{R}^{D},\\ &\ (\boldsymbol{\theta},{\bf y})\mapsto{\bf f}(\boldsymbol{\theta},{\bf y})\end{split}$
such that
${\bf f}(\boldsymbol{\omega}t,{\bf y})\equiv e^{-t\mathcal{A}}{\bf g}(e^{t\mathcal{A}}{\bf y}).$
Therefore, the change of variables (42) transforms (1) into the quasi-periodic system
$\frac{d}{dt}{\bf y}={\bf f}(\boldsymbol{\omega}t,{\bf y}).$ (43)
The definition of semi-discrete Taylor-Fourier approximations follows naturally: in Subsection 2.1, replace the basic frequency $\omega$ and the angular variable $\theta$ by the vector of non-resonant frequencies $\boldsymbol{\omega}$ and the angular variables $\boldsymbol{\theta}$, respectively. The discussion on high order stroboscopic averaging in Section 4 is also valid in the quasi-periodic case with the referred changes (the averaging technique is referred in that case as quasi-stroboscopic averaging [4, 5]).
Fully discrete $(M,d)$-Taylor-Fourier approximations extend this approach by employing trigonometric interpolation of functions $\mathbb{T}^{r}\to\mathbb{C}$ over an $r$-fold Cartesian product of the uniform grid
$\theta_{n}=\frac{n\pi}{M},\quad n=0,1,2,\ldots,2M-1,$
in $[0,2\pi)$, which can be efficiently computed using multi-dimensional discrete Fourier transforms (DFT).
However, compared to the periodic case, quasi-periodicity requires increased memory and higher computational cost: while a complex tensor of dimensions $D\times(d+1)\times(M+1)$ suffices to store the coefficients ${\bf y}^{[M,d]}_{k,j}$ in the periodic case, the quasi-periodic setting with $\boldsymbol{\omega}\in\mathbb{R}^{r}$ requires a complex tensor of dimensions $D\times(d+1)\times(M+1)^{r}$, significantly increasing storage and processing demands.
5.3 Equations in other alternative formats
Going back to equations in the real domain, Taylor-Fourier approximations are not limited to semi-linear equations of the form (2). They can be applied more generally to problems of the form
  1. 1.
    $\frac{\partial}{\partial t}\varphi_{t}({\bf x})={\bf r}(\varphi_{t}({\bf x})).$ (45)
    In that case, problem (44) is transformed with the change of variables
    ${\bf x}(t)=\varphi_{\omega t}({\bf y}(t))$
    into (5), where
  2. 2.
${\bf x}^{[M,d]}(t)=\varphi_{\omega\,t}\left({\bf y}^{[M,d]}(t)\right),$ (47)
where
${\bf y}^{[M,d]}(t)=\sum_{k=-M}^{M}e^{ik\omega t}\sum_{j=0}^{d}t^{j}\,{\bf y}_{k,j}^{[M,d]}$
is the $(M,d)$-Taylor-Fourier approximation to the solution of (5) with the map ${\bf f}$ defined by (46).
$\begin{split}\frac{d}{dt}\phi_{j}&=k_{j}\omega+g_{j}({\bf x}),\quad j=1,\ldots,s,\\ \frac{d}{dt}a_{j}&=g_{s+j}({\bf x}),\quad j=1,\ldots,D-s,\\ \end{split}$ (48)
where ${\bf x}=(\phi_{1},\ldots,\phi_{s},a_{1},\ldots,a_{D-s})$, $k_{j}\in\mathbb{Z}$, and for each $j$, $g_{j}$ is $2\pi$-periodic in each of the angle variables $\phi_{j}$. In that case,
$\varphi_{t}(\phi_{1},\ldots,\phi_{s},a_{1},\ldots,a_{D-s})=\left(\begin{matrix}\phi_{1}+k_{1}\,t\\ \vdots\\ \phi_{s}+k_{s}\,t\\ a_{1}\\ \vdots\\ a_{D-s}\\ \end{matrix}\right),$
and the change of variables ${\bf x}(t)=\varphi_{\omega t}({\bf y}(t))$, with ${\bf y}=(\tilde{\phi}_{1},\ldots,\tilde{\phi}_{s},\tilde{a}_{1},\ldots,\tilde{a}_{D-s})$, transforms the system (48) into
$\frac{d}{dt}\tilde{\phi}_{j}$ $=g_{j}(\tilde{\phi}_{1}+k_{1}\,\omega\,t,\ldots,\tilde{\phi}_{s}+k_{s}\,\omega\,t,\tilde{a}_{1},\ldots,\tilde{a}_{D-s}),\quad j=1,\ldots,s,$
$\frac{d}{dt}\tilde{a}_{j}$ $=g_{s+j}(\tilde{\phi}_{1}+k_{1}\,\omega\,t,\ldots,\tilde{\phi}_{s}+k_{s}\,\omega\,t,\tilde{a}_{1},\ldots,\tilde{a}_{D-s}),\quad j=1,\ldots,D-s.$
@@ 6 Numerical illustrations
In this section we include two numerical examples to illustrate the performance of the Taylor-Fourier approximations proposed in this paper. All codes we have used to carry out the numerical experiments are available in https://github.com/jmakazaga/Taylor-Fourier.
@@ 6.1 The cubic nonlinear Schrödinger equation
Products of a vector in $\mathbb{C}^{2J}$ by matrices $e^{\theta A}$ and $e^{-\theta A}$ can be computed efficiently by performing two FFT in $\mathbb{C}^{2J}$, since $e^{\theta A}={\cal F}^{-1}e^{\theta D}{\cal F}$, and $e^{\theta D}$ is the diagonal matrix with diagonal entries $(e^{\lambda_{1}\theta},\ldots,e^{\lambda_{2J}\theta})$. On the other hand, due to the particular form of the nonlinearity ${\bf g}$ in (53), the Taylor expansions required to compute the Taylor-Fourier approximations are cheaply available since sums and products of power series in $t$ can be computed in a very efficient way.
@@
Secondly, to illustrate how the choice of $d$ and $M$ affects the approximate solution computed with the Taylor-Fourier method, we include in Figure 2 the numerical approximations to the solution of (49)-(55) at $t=\pi/10$ when $J=2^{9}$, and the combinations $M=4J$, $d=1$ (left), $M=J$, $d=3$ (center) and $M=J$, $d=1$ (right) are used. Plots in Figure 2 must be compared with the bottom right plot in Figure 1 where the $(4J,3)$-Taylor-Fourier approximation is displayed. Comparing the plots corresponding to the same value of $M$ and reducing $d$ from 3 to 1, we observe that, although the overall shape of the numerical solutions is rather similar, the maximum and minimum values of both solutions are clearly different. Looking now at the solutions obtained with the same value of $d$ we see that when decreasing $M$ from $4J$ to $J$ small oscillations appear.
@@ 6.2.2 Stiefel and Scheifele’s formulation
${\bf u}(\tau)=\cos(\omega\tau)\boldsymbol{\alpha}(\tau)+\omega^{-1}\sin(\omega\tau)\boldsymbol{\beta}(\tau),$
where $\boldsymbol{\alpha}(\tau)$ and $\boldsymbol{\beta}(\tau)$ are the solutions of the 8-dimensional ODE system
$\begin{split}\boldsymbol{\alpha}^{\prime}&=\omega^{-1}\sin(\omega\tau)\nabla R(\cos(\omega\tau)\boldsymbol{\alpha}+\omega^{-1}\sin(\omega\tau)\boldsymbol{\beta}),\\ \boldsymbol{\beta}^{\prime}&=-\cos(\omega\tau)\nabla R(\cos(\omega\tau)\boldsymbol{\alpha}+\omega^{-1}\sin(\omega\tau)\boldsymbol{\beta}),\end{split}$ (67)
@@
$\boldsymbol{\alpha}(0)={\bf u}_{0},\quad\boldsymbol{\beta}(0)={\bf u}^{\prime}_{0}.$ (68)
@@
$\boldsymbol{\alpha}(\tau)\approx\sum_{k=-M}^{M}\sum_{j=0}^{d}e^{ik\omega\tau}\,\tau^{j}\,\boldsymbol{\alpha}_{k,j},\quad\boldsymbol{\beta}(\tau)\approx\sum_{k=-M}^{M}\sum_{j=0}^{d}e^{ik\omega\tau}\,\tau^{j}\,\boldsymbol{\beta}_{k,j},$
of $\boldsymbol{\alpha}(\tau)$ and $\boldsymbol{\beta}(\tau)$ are obtained.